### v1

# Chatter: A Character Attribution Dataset for Narrative Understanding

Abstract
Computational narrative understanding studies the identification, description, and interaction of the elements of a narrative: characters, attributes, events, and relations. Narrative research has given considerable attention to defining and classifying character types. However, these character-type taxonomies do not generalize well because they are small, too simple, or specific to a domain. We require robust and reliable benchmarks to test whether narrative models truly understand the nuances of the character’s development in the story. Our work addresses this by curating the Chatter dataset that labels whether a character portrays some attribute for 88148 character-attribute pairs, encompassing 2998 characters, 13324 attributes and 660 movies. We validate a subset of Chatter, called ChatterEval, using human annotations to serve as an evaluation benchmark for the character attribution task in movie scripts. ChatterEval assesses narrative understanding and the long-context modeling capacity of language models.
Chatter: A Character Attribution Dataset for Narrative Understanding
Sabyasachee Baruah1  and Shrikanth Narayanan2 Signal Analysis & Interpretation Laboratory University of Southern California sbaruah@usc.edu1, shri@usc.edu2
1 Introduction
Narrativity occurs when characters interact with each other, triggering events that are temporally, spatially and causally connected. This sequence of events forms the story. Piper et al. (2021) provided a symbolic definition of narrativity wherein they asserted that narrativity occurs when narrator $\mathcal{A}$ tells perceiver $\mathcal{B}$ that some agent $\mathcal{C}$ performed the action $\mathcal{D}$ on another agent $\mathcal{E}$ at place $\mathcal{F}$ and time $\mathcal{G}$ for some reason $\mathcal{H}$. Baruah and Narayanan (2024) used this definition to identify four main elements of any narrative: characters, attributes, events and relations. Labatut and Bost (2019) explored different types of character interactions and emphasized the central role characters play in narratives. Characters drive the plot forward through their actions, develop attributes, arouse tension and emotion in the story by creating conflicts or bonds with other characters, and embody tropes and stereotypes to relate to the audience. The vitality of characters in narratives makes character understanding an essential task in narrative research.
Dataset Domain Type Size
Finlayson (2015) Folklore Archetype 282
Skowron et al. (2016) Movies Role 2010
Brahman et al. (2021) Literature Description 9499
Sang et al. (2022) Movies Personality 28653
Yu et al. (2023) Literature Traits 52002
Chatter Movies Tropes 88148
ChatterEval Movies Tropes 1062
Table 1: Comparison with other attribution datasets. Size denotes the number of character-attribute pairs.
Narratologists have explored various approaches to operationalize the character understanding task. For example, Inoue et al. (2022) presented character understanding as a suite of document-level tasks that included gender and role identification of the character, cloze tasks, quote attribution and question answering. Li et al. (2023) adopted coreference resolution, character linking and speaker guessing tasks, and Azab et al. (2019) used character relationships and relatedness to evaluate character representations. We organized the character understanding tasks into the following categories. 1) Identification tasks find the unique set of characters and their mentions. It includes entity recognition, entity linking and coreference resolution. The 2) Quotation task maps utterances to characters. 3) Attribution tasks, such as personality classification, persona modeling and description generation, describe the character. The 4) Cloze task asks the model to fill in the correct character name given an anonymized character description, story summary or story excerpt. 5) Relation tasks, such as relation classification, draw similarities between characters.
Among these tasks, character attribution is the most challenging because there exists a multitude of ways to qualify a character, such as personality (Sang et al., 2022), adjectives (Yu et al., 2023), persona (Bamman et al., 2013, 2014), archetypes (Finlayson, 2015), roles (Skowron et al., 2016) and descriptions (Brahman et al., 2021). Each of these approaches has drawbacks. For example, personality scales such as Big5 and MBTI cannot capture all the variation in characterization and can be difficult to interpret (Zillig et al., 2002). Adjectival descriptors are too general and only apply to a limited context in the story. Persona roles are not explicitly defined. Archetypes (Propp, 1968; Jung, 2014) are abstract concepts specific to the domain of interest. Character descriptions are detailed, freeform and scalable to any number of characters. However, we cannot factor them into simpler components or use them to compare between different characters.
We require a robust attribution taxonomy that can scale across different narratives, characters and domains, is well-defined and discrete for effective character comparison, and necessitates document-level understanding to model them accurately. We developed the Chatter dataset to fulfill this need. We used character tropes from the TVTropes website as our attribute types, and drew our characters primarily from Hollywood movies across various genres. Chatter labels indicate whether the character portrayed the trope in the movies they appeared in, encompassing 88148 character-trope pairs. It also provides the full-length screenplays of the movies, averaging about 25K words per screenplay. We define the character attribution task as a binary classification task where, given a character-trope pair and the screenplays of the movies where the character appears, the model should predict whether the character portrayed the trope. We validate a subset of the Chatter data, referred to as ChatterEval, using human annotations to establish an evaluation benchmark for the character attribution task. We compared the zero-shot performance of LLMs and Chatter’s labels to assess the suitability of Chatter as a training set for the attribution task. The dataset is available at https://drive.google.com/drive/folders/11egMhs-zkWSASe7zJENwHg17-6VOeXDU?usp=sharing.
2 Data
2.1 Tropes
Tropes are storytelling devices or conventions used by the writer to easily convey some story notion to the reader. They act like narrative motifs that readers can easily recognize, saving time and effort for the writer as they can omit details the readers can infer from the trope. For example, the AntiHero is a very popular trope that describes a protagonist who lacks traditional heroic qualities, often cynical and flawed, yet ultimately performs heroic actions. The character Severus Snape in the Harry Potter stories portrays the AntiHero trope as he gives Harry a hard time but stays loyal to Dumbledore (mentor) and works secretly to defeat Voldemort (antagonist).
Tropes can also relate to inanimate objects, events, locations or the environment. We focus only on character tropes in our work. The scope of a trope can vary greatly. Some tropes like PetTheDog (villain performing an act of kindness) are portrayed over short contexts, typically a single scene, whereas others like HiddenDepths (revealing unexpected talents as the story progresses) are portrayed over a longer context. The attribution model should be able to extract information from different points in the narrative and reason over it to identify the portrayed tropes.
We use the character trope labels of TVTropes [^1] https://tvtropes.org. TVTropes is a community-driven website, similar to Wikipedia, that catalogs tropes with definitions and examples. Fans of a creative work discuss and post tropes they identify in the narrative. The character trope annotations we collect from TVTropes are those perceived by the reader and not the actual tropes intended by the creator. Since there is no quality control or quantifiable agreement on the published content, we treat the TVTropes data as a noisy source of character attribution.
2.2 Screenplays
We used movies as the source of our narratives. We chose the cinematic over the literary domain because it had more TVTropes labels, and we supposed it would be easier to find raters knowledgeable about movies than books. Additionally, movies allow us to extend the attribution task to the multimodal domain, thereby extending its scope.
We used publicly available movie screenplays from the ScriptsonScreen [^2] https://scripts-onscreen.com website. Each script is mapped to an IMDB [^3] https://imdb.com identifier so we can uniquely identify the movie. Most movies in our dataset are produced in the US or the UK after 1980. The average script size is about 25K words. We apply a named entity classifier and name alias generator to map the character names in the script to a unique character in the IMDB cast list. We preprocess the screenplays to find scenes, dialogues and descriptions using Baruah and Narayanan’s (2023) screenplay parser. In total, our dataset contains screenplays of 660 movies.
2.3 Chatter
We build the Chatter dataset of character-trope pairs using the tropes of TVTropes and the screenplays of ScriptsonScreen. First, we download the movie screenplays from ScriptsonScreen, parse and map them to an IMDB page, and match the characters occurring in the document to a character in the IMDB cast list. Second, we search for the character in TVTropes and retrieve their character page. Third, we collect the tropes portrayed by the character from the character page, as labeled by the TVTropes community. Fourth, we search for the trope page and fetch its definition. We also summarize the definition using GPT-4 for the annotation task. The average size of the trope definition and its summary is 344 and 44 words, respectively.
We need good negative samples to evaluate the specificity of the attribution model. However, TVTropes does not provide this information because it does not label tropes not portrayed by the character. Instead, we analyze the trope definition to find antonym tropes and create the negative character-trope pairs. For example, the definition of the AntiHero trope contains the sentence – "…Compare and contrast this trope with its antithesis, the AntiVillain…" – which indicates that the AntiVillain trope is contrary to AntiHero. We search the trope definitions to retrieve opposing tropes by checking if negation words such as "contrast", "opposite" and "counterpart" exist within a five-word context window of the antonym trope mention.  [^4] The full list of negation words is included in the dataset For each positive character-trope pair, we create a negative pair by either – 1) selecting an antonym trope from the trope definition or 2) choosing a trope that is absent in any of the positive character-trope pairs for that character. We randomly choose between the two methods with 50% probability to get a good mix of hard and soft negative samples.
We add the negative samples to complete the construction of Chatter. It contains 88148 character-trope pairs with a 50-50 split between positive and negative pairs, encompassing 2998 unique characters from 660 movies and 13324 tropes.
3 Evaluation Data
We can use Chatter to train attribution models, but we require a more reliable dataset for evaluation. We annotate a subset of Chatter using human raters and create the ChatterEval dataset.
3.1 Annotation
We sampled movies from our dataset released after 2010 and collected the corresponding character-trope pairs for annotation. We employed workers on the Amazon MTurk [^5] https://www.mturk.com crowdsource annotation website. We selected workers with high reputation and experience666 ¿98% approval rate and worked on ¿5000 HITs and tested them on two separate qualification tasks, each containing five questions that asked them to decide whether the character portrayed the given trope. The task showed them the picture of the character, the movies that starred the character with Wikipedia links, a link to the fandom page [^7] https://www.fandom.com where the worker can read more about the character, and the summarized trope definition with a link to the trope’s TVTropes page. We also surveyed the workers about the movies they had watched. We qualified 69 workers who correctly answered at least 80% of the questions.
The qualified workers labeled whether the character portrayed the trope on a Likert scale: no, maybe no, not sure, maybe yes and yes. To reduce the cognitive load on the workers, they annotated the character-trope pairs in batches, each batch containing characters from at most ten movies. Before each batch, we inform the workers about the movies to prepare them for the annotation task. After each batch, we estimate the worker’s performance by calculating the accuracy of their ratings against the Chatter labels and record how much time they spent on the task. We dropped workers whose accuracy dropped below 50% or those speeding through the samples, disqualifying them from working on future batches. Of the 69 workers, we disqualified 10 raters and were left with 59 reliable workers.
The annotation task ran for one month. We paid the workers at the rate of \$18/hour, and the entire task cost us about \$10K. We collected three annotations per character-trope pair, totaling 5683 annotations. The Krippendorff inter-rater reliability score is 0.448.
Dataset Characters Tropes Movies Samples
Chatter 2998 13324 660 88148
ChatterEval 271 897 78 1062
Table 2: Chatter and ChatterEval’s data statistics.
3.2 ChatterEval
We aggregate the annotations to build the ChatterEval dataset. The character-trope pairs do not all show clear agreement. Character attribution, like sentiment analysis, is a subjective task, and whether a character portrays some trope is perceived differently by people. We drop the very ambiguous samples and those with insufficient reliable ratings (refer to Appendix A). We assign a confidence score between 0 and 1 to the remaining samples based on the agreement between the raters. Models can use these during evaluation to weigh the mispredictions.
In total, ChatterEval contains 1062 labeled character-trope pairs, covering 271 characters, 897 tropes and 78 movies. It contains 556 positive and 506 negative pairs. We include the annotations of the individual raters in ChatterEval to encourage multiannotator modeling. Table 2 shows some data statistics of Chatter and ChatterEval.
4 Experiments
4.1 Models
We establish baselines on ChatterEval using zero-shot prompting. We use Gemini-1.5-Flash (Reid et al., 2024) and Llama-3.1-8B-Instruct (Dubey et al., 2024) because they can handle long contexts. We experiment with two input settings. We include the full screenplay in the prompt or only provide the segments where the character speaks or is mentioned. The prompt then contains the definition of the trope and asks whether the character portrays the trope. We set the temperature to 1 and average the performance over five runs. We also evaluate Chatter’s labels against the annotations of ChatterEval to assess its suitability as a training set for character attribution.
4.2 Results
Table 3 shows the performance of the baseline models. Prompting the character-specific segments significantly improves performance over using the full script because the model does not have to search the relevant information in the long document. The superior performance of Gemini-1.5 over Llama-8B is expected because it has more parameters. Prompting Llama-8B on the full script performed even worse than random guessing, showing that there is much scope for improvement in locally trained attribution models. Chatter can serve as a good training dataset for that purpose because its labels showed the best performance on the test set. We also observed that the models consistently predicted that the character did not portray the trope (low recall). The lack of information due to intractable pairs or the limitation of the text modality could be contributing to the high false negatives.
Input Model Precision Recall F1 w-F1
Random 52.3 50.0 51.1 51.1
Chatter’s labels 81.0 82.2 81.6 83.6
Script Llama-8B 61.9 24.8 35.4 36.1
Gemini-1.5 92.3 39.7 55.5 59.1
Segment Llama-8B 62.9 60.5 61.7 62.8
Gemini-1.5 88.1 66.3 75.7 79.3
Table 3: Zero-shot performance on ChatterEval for the character attribution task. w-F1 is the weighted F1 score. Chatter denotes Chatter’s labels.
5 Related Work
Several past studies have curated character attribution datasets for narrative understanding. Finlayson (2015) annotated seven character archetypes in Russian folktales to create the ProppLearner corpus. Skowron et al. (2016) labeled hero, antagonist, sidekick, spouse and supporting roles in action-genre movies. Brahman et al. (2021) collected character descriptions from online study guides such as LitChart and Sparknotes, and created the LiSCU dataset for the character identification and description generation task. Sang et al. (2022) curated MBTI personality types from the personality database for movie characters. Yu et al. (2023) annotated reader’s notes for character traits in Chinese-translated Gutenberg books. Baruah and Narayanan (2024) annotated screenplay excerpts for character attributes and evaluated in-context and chain-of-thought learning methods on the attribution task. Table 1 compares these datasets against Chatter and ChatterEval.
6 Conclusion
We proposed the Chatter and ChatterEval datasets for the narrative character attribution task. We addressed the limitations of previous datasets by curating a resource that is scalable, generalizable, well-defined and discrete. Experiments showed that Chatter can serve as a reliable training set and ChatterEval can be used as the evaluation benchmark for character attribution modeling. Future work includes developing character attribution models on our datasets to aid creators and writers in analyzing their narratives.
7 Limitations
We define the character attribution task as a binary classification task, where, given the character-trope pair and the screenplays of the movies in which the character appeared, the model should predict whether or not the character portrayed the trope. This formulation has some limitations. First, the screenplay’s narrative may not exactly resemble the story told by the movie because of tweaks made during the filming process. Publicly available screenplays are rarely the final script but drafts from earlier in the production stage. Second, movies could have visual cues like the non-verbal behavior of the character that are missed by the screenplay text. Lastly, a character can appear in multiple movies, and our dataset does not contain scripts for all of them. These limitations have important implications because the TVTropes contributors and our raters draw their knowledge of the character from all sources. In contrast, the attribution model predicts the attribute labels based only on the screenplay. There could be character-attribute pairs for which the model has insufficient information and could lead to poor recall, as shown by the performance of the baseline models in Table 3.
Appendix A Annotation
We give the following annotation instructions to the Amazon MTurk workers.
A trope is a storytelling device or
convention the storyteller uses to
describe situations the audience can
easily recognize, often used to
stereotype characters. Think of them
as character attributes.
Your task is as follows:
1. Identify the movie character from the
   pictures and the given movie(s). If
   you are having trouble recognizing
   the character, click on the Character
   Page link(s).
2. Read the definition of the character
   trope. You can open the link to know
   more.
3. Choose yes, maybe yes, not sure,
   maybe no, or no to answer whether the
   character portrays or is associated with
   the trope.
4. (Optional) Explain your choice in the
   provided text area or leave any other
   comment behind.
Figure 1: Interface of the annotation task.
Figure 1 shows the annotation interface that the workers use to label character attribution. The annotation task ran from August 11 to September 19, 2024. Throughout the task, we maintained a communication channel for the workers in the TurkerNation Slack Workspace, where we responded to any questions the workers had about the task. We pay \$1.5 for each question which turns out to be \$18/hour because the workers spent an average of 5 minutes per question.
The MTurk workers answer yes, maybe yes, not sure, maybe no or no on each question. We map this ordinal scale to a numeric range by mapping the labels to 2, 1, 0, -1 and -2. Each character-trope pair is annotated by at most three workers. We sum the label values for each sample to get an integer score between -6 and 6. The higher the absolute score, the more agreement between the raters. We drop samples whose absolute integer score was below 3. For the remaining samples, we obtain the annotation confidence $w$ by normalizing the integer score $s$ between 0 and 1: $w=(s-2)/4$. The confidence score takes values 0.25, 0.5, 0.75 and 1. Attribution models can use these scores to weigh their evaluation metrics. Table 3 shows the weighted F1 scores for the baseline models.
Appendix B TVTropes
TVTropes is licensed under a Creative Commons Attribution-NonCommercial-ShareAlike 3.0 Unported License. This license allows for the copy and redistribution of the material in any medium or format for non-commercial purposes. Chatter is published for academic research purposes and does not infringe the TVTropes license.
Appendix C ScriptsonScreen
ScriptsonScreen aggregates movie scripts from different websites such as IMSDB [^8] https:https://imsdb.com/, Dailyscripts [^9] https://www.dailyscript.com/ and Script Slug [^10] https://www.scriptslug.com/. All the movie scripts we use are licensed for fair use and available for public download.
Appendix D Prompt
We used the following prompt template in our experiments.
Given below is a movie script enclosed
between the tags <SCRIPT> and </SCRIPT>.
Following that is the definition of the
\$\$TROPE\$\$ trope enclosed between the tags
<TROPE> and </TROPE>.
Read the movie script and based on that
answer yes or no if the character
\$\$CHARACTER\$\$ portrays the \$\$TROPE\$\$
trope.
Also, give a brief explanation as to why
you think the answer is yes or no.
The explanation should follow after the
yes/no answer.
<SCRIPT>
\$\$SCRIPT\$\$
</SCRIPT>
<TROPE>
\$\$DEFINITION\$\$
</TROPE>
We replace \$\$CHARACTER\$\$, \$\$TROPE\$\$, \$\$DEFINITION\$\$ and \$\$SCRIPT\$\$ with the character name, trope name, definition of the trope and the movie script during prompting. We do not ask for an explanation when prompting Llama-3.1-8B-Instruct.
Appendix E Experiments
We used the closed-source Gemini-1.5-Flash and the open-source Llama-3.1-8B-Instruct models to develop the baselines. We used the Gemini API to access Gemini-1.5-Flash. We ran the inference pipeline of Llama-3.1-8B locally on one NVIDIA A40 GPU. The numbers reported in Table 3 were averaged over five runs.